OpenStudy (bluefire113):

Simplify e^4x-2e^x+1 Thank you!

jhonyy9 (jhonyy9):

note e^x = y and will get y^2 -2y +1 or 1 = e^0 but just you need simplify it or solve it for x ?

OpenStudy (bluefire113):

just simplify it

jhonyy9 (jhonyy9):

than just simplify you can factorizing out the e^x e^x(e^3x -2) +1

OpenStudy (bluefire113):

would the answer be e^5x-1?

zepdrix (zepdrix):

As jhonyy metioned, making a substitution of sorts gives us,\[\large\rm y^2-2y+1\]This is a perfect square trinomial, do you know to factor it?

OpenStudy (bluefire113):

yes that would be (y-1)^2

zepdrix (zepdrix):

\[\large\rm \left(\color{orangered}{y}-1\right)^2\]Good. Then undo your substitution you made,\[\large\rm \left(\color{orangered}{e^x}-1\right)^2\]

OpenStudy (bluefire113):

okaay..

satellite73 (satellite73):

in what way is factoring the same as "simplifying" ? answer, in no way ask your math teacher what he/she means by "simplify"

OpenStudy (bluefire113):

uh, the question says simplify.

OpenStudy (bluefire113):

i think you have to multiply the two e's.

OpenStudy (bluefire113):

the equation looks like this: (e^4x-2)*(e^x+1)

satellite73 (satellite73):

it is really \[e^{4x}-2e^x+1\]?

satellite73 (satellite73):

hold the phone are you suppose to multiply \[(e^{4x}-2)\times (e^x+1)\] is that it?

OpenStudy (bluefire113):

not really

satellite73 (satellite73):

what was the actual question ?

OpenStudy (bluefire113):

ohh....i think i know why you are confused, the 4x-2 is part of the exponent as well as the x+1

OpenStudy (bluefire113):

so its: e^(4x-2)e^(x+1)

satellite73 (satellite73):

ooh \[\huge e^{4x-2}\times e^{x+1}\]??

OpenStudy (bluefire113):

YES!!

satellite73 (satellite73):

add the exponents just like with any base

satellite73 (satellite73):

i.e. add \[4x-2+x+1\] by combining like terms, stick that up in the exponent

OpenStudy (bluefire113):

oh ok so the answer is e^(5x-1)

satellite73 (satellite73):

yes

OpenStudy (bluefire113):

:) Thank you very much and sorry for the confusion!